none of my answers are working, please help. Chapter 6, Section 2-CI, Exercise 082 Your answer is partially correct. Try again. Use the t-distribution to find a confidenc
Answer #1

Here, given that, X= 75.3,5= 8.0, n=42 point estimate = 175.3 margin of error tn-1;4/2 2 1 where, tnijelz= tez-1; 4/2= taljooHi Dear,
I have given my 100% to solve your problem.
So, please help me out by just thumbs up.
Thank you, so much.